Find the whole quantity if
 5% of it is 600
Find the whole quantity if 
 5% of it is 600

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us assume the whole quantity be x, Then,
(5/100) × (x) = 600
X = 600 × (100/5)
X = 60000/5
X = 12000

Thanks